OpenStudy (anonymous):

Which of the following could be the lengths of the three sides of a triangle? A.16, 18, 30 B.5, 5, 11 C.20, 40, 60 D.6, 6, 12

OpenStudy (anonymous):

use pythagoras theorem

ganeshie8 (ganeshie8):

To form a triangle : sum of any two sides must be greater than the third side

ganeshie8 (ganeshie8):

consider option D : D.6, 6, 12 since 6+6 is not greater than 12, these lengths cannot form a triangle

ganeshie8 (ganeshie8):

see if you can check other options similarly

OpenStudy (anonymous):

It needs to know what 3 sides could make a triangle .

OpenStudy (anonymous):

for making a triangle tthe sum of two sides should be equal to or greater than the third side

OpenStudy (anonymous):

i know but i cant figure it out, but thank you for the help (:

OpenStudy (anonymous):

wlcm

OpenStudy (gorv):

like option c 20, 40 ,60 take two side 20, 40 3rd side = 60 sum = 20+40=60 20, 60 3rd side =40 sum =20+60=80 40, 60 3rd side = 20 sum = 40+60=100 in any of 3 combination if atleast one combination is their sum of ehose 2 side > 3rd side than it can form triangle ..similarly try for others